### v1

# Division of labor enables efficient collective decision-making under uncertainty

Abstract
How do social animals make effective decisions in the absence of a leader? While coordination can improve accuracy, it also delays responses as information propagates through the group. In changing environments, these delays can outweigh the benefits of centralized control, making decentralized strategies advantageous in large groups. This raises a key question: how can groups implement efficient collective decisions without central coordination? We address this question using a model of collective foraging in which individuals choose whether to invest in costly exploration or remain idle, while sharing information and rewards across the group. We show that decentralized collectives can match the performance of centrally controlled groups through a division of labor: a small, heterogeneous subset explores even when expected rewards are negative, while a synchronized majority forages only when expected rewards are positive. Information redundancy causes the optimal scout number to grow sublinearly with group size, so that larger groups need proportionally fewer explorers. The heterogeneity of the group is maximized at intermediate ecological pressures, but optimal groups are homogeneous when costs or environmental contrasts or fluctuations are extreme. Crucially, these group-level policies do not require central coordination, emerging instead from agents following simple threshold-based decision rules. We thus demonstrate a mechanism through which leaderless collectives can make effective decisions under uncertainty and show how ecological pressures can drive changes in the distribution of strategies employed by the group.
1 Introduction
How animal groups balance individual exploration costs and collective information gains remains a central question in understanding ecological decision-making [Conradt2003, Couzin2005, Sumpter2005, Couzin2009]. Individual foragers bear energetic and survival costs when venturing into uncertain environments [Stephens1986, Wolf1989], yet the information about the environmental state resulting from these forays can outweigh any immediate rewards [Couzin2005, Hein2015, Dall2005]. This tension raises a fundamental question: How should groups be organized to balance these costs and benefits?
One solution to this problem is centralized leadership. In many wild wolf packs the breeding pair tends to initiate movements and coordinates traveling, resting, and hunting [Mech1999]. Human history offers parallel examples. Admiral Zheng He commanded fleets across the Indian Ocean in the early fifteenth century on imperial orders, spurring Chinese emigration and trade throughout Southeast Asia [Levathes2014]. A century later, royal sponsorship supported Magellan’s circumnavigation, which opened the first western sea route to the Spice Islands and accelerated global trade and contact between distant regions [Bergreen2003]. In each case, a designated explorer, acting under a central authority, bore enormous personal risk, while the nations that sent them reaped lasting rewards. However, most social animals including honeybees, ants, and fish schools lack fixed leaders and coordinate through local interactions [Pratt2002, Seeley2004, Couzin2005, Sumpter2005, Seeley2011, Hein2015]. Despite extensive empirical study of decentralized systems [Pratt2002, Sumpter2005, Giardina2008, Gordon2014], it remains unclear how closely decentralized coordination matches centralized control, how performance depends on diversity in individual strategies, and under what conditions heterogeneous strategies offer an advantage in leaderless groups.
To address these questions we develop and analyze a stochastic foraging model [Puterman1994, Mahadevan1996, Gold2007, Kilpatrick2021, BlumMoyse2025, Dembo2010, Bressloff2021, Barendregt2025] to understand how collectives can achieve efficient exploration. We show that well-calibrated collectives exhibit characteristic dynamics: during times of plenty, risk-averse individuals venture out together, amplifying collective gains. When resources are scarce, most individuals remain idle. Yet, a small subset of risk-tolerant explorers continues foraging even when expected rewards are negative, and can rapidly detect environmental improvements. Composition is critical as population dominated by cautious individuals fail to discover new resources, while those dominated by risk-tolerant foragers exhaust shared reserves. Indeed, the number of risk-tolerant individuals grows sublinearly with population size, making a balanced division of labor between exploration and exploitation essential for efficient foraging [Seeley2011, Charbonneau2015]. Remarkably, collectives with a well-calibrated mixture of cautious and bold individuals can perform as well as groups governed by a central decision-maker. The most heterogeneous compositions emerge under intermediate ecological pressures, when exploration costs and environmental variability are balanced. We thus show how simple individual decision rules lead to optimal collective performance without coordination, and predict when ecological pressures favor heterogeneous versus uniform group composition.
Figure 1: Overview of model and setup. Individuals in a collective observe returning foragers to update their belief about the environmental state and select actions. (A) The belief, $g^{t}$, is updated based on the observed outcomes of foraging attempts. (B) Individuals forage when their belief that the environment is in a high-reward state exceeds a threshold, $\theta$; these thresholds can differ between individuals. (C) Foraging is successful with probability $\gamma^{t}$, which switches between high and low values with the environmental state. (D) Under centralized control, a planner decides on the optimal number of foragers to maximize returns. (E) Under decentralized control, coordination emerges from individual foraging decisions.
2 A Model of Collective Foraging under Uncertainty
Collective decisions about foraging in large, decentralized collectives face a fundamental challenge: How do groups balance individual risk and collective benefit in dynamically changing environments without central coordination? To address this question we consider a minimal model that shows that decentralized decision-making mechanisms can yield optimal collective long-term returns, even when individual expected returns are negative. Here we provide an overview of the model; mathematical details can be found in SI Appendix, section 1.
Foraging agents in a dynamic environment
We consider a collective of $N$ agents who forage for resources. Time is discrete, and on each time step an agent chooses either to leave the “hive” and forage, or to stay. Each foraging attempt is either successful, resulting in a positive reward, or unsuccessful, resulting in no reward. The environment alternates between a high- and low-rewarding state, so that in the high (low) state a foraging attempt is successful with probability $\gamma_{+}$ ($\gamma_{-}$). To model energetic expenditure and environmental risk, we assume that each foraging attempt incurs a fixed cost, $\lambda$, and that the expected return from an attempt is positive in the high-reward state and negative in the low-reward state. Transitions between the states are memoryless, so the environmental state evolves as a two-state discrete-time Markov process with symmetric switching probability, $\epsilon$ (Fig. 1C). While real environments fluctuate across a continuum of states, assuming two states provides a minimal framework that captures the essential trade-off between exploration and exploitation under uncertainty [Gold2007, Kilpatrick2021, Barendregt2025, BlumMoyse2025].
Collective perception
We assume that the community coordinates its behavior through shared information about the outcomes of foraging attempts. Returning foragers are observed by the entire community, so all individuals know how many foraging attempts have been successful (Fig. 1A). Individuals are Bayesian observers [Wald1948, Bogacz2006, VelizCuba2016] and compute the probability that the environmental state is in a high-reward state based on the observed number of successful foragers and knowledge about the environment’s volatility. We refer to this subjective probability as their belief about the environmental state. Because all individuals have access to the same information, they share a common belief that guides their foraging decisions.
Decision-making with a centralized planner
Maximizing returns requires optimizing forager number: too many foragers when conditions are poor wastes effort; too few when conditions are good misses opportunities. To establish an upper bound on collective performance, we derive a theoretical benchmark using a centralized planner that optimally determines foraging numbers to maximize long-term collective returns (Fig. 1D). Under this centralized policy, all individuals follow the directives of a planner. This planner infers the environmental state by observing foraging successes to determine the best number of foragers. This optimal policy can be computed using dynamic programming [Bellman1954, Mahadevan1996, Sutton1998] and provides an upper bound on achievable performance. We compare this benchmark with outcomes from independent agents who make individual decisions based on shared information but without central coordination.
Decisions in the absence of a centralized planner
With no central decision-maker, each individual decides whether to forage based on their estimate of the environmental state. We assume that each agent chooses to forage if their belief that the environment is in the high-reward state exceeds a private threshold (Fig. 1B, E). The distribution of these thresholds across individuals determines the behavior of the decentralized group [Bonabeau1996, Bonabeau1998]. Such threshold-based decision rules are consistent with empirical work on quorum/acceptability thresholds in social insects and with theory showing how heterogeneity in thresholds shapes collective decisions [Theraulaz1998, Franks2003, Robinson2011, Conradt2012, Jeanson2014, Masuda2015].
Despite its abstraction, this model captures the ecological tradeoffs between exploration cost, environmental uncertainty, and collective return. It also allows us to ask how well a decentralized collective can perform compared to an optimally controlled group.
3 Results
We use our model of foraging and belief-updating to ask how a collective translates shared information about the environment into an effective foraging decisions. We begin by establishing a normative benchmark defined by the allocation selection of a centralized decision-maker to maximize expected payoff. We then show that a leaderless group can achieve optimal performance using heterogeneous individual threshold rules, and analyze how environmental conditions shape structured heterogeneity in an optimal, decentralized collective.
3.1 Robust collective coordination through structured heterogeneity
A central decision-maker can select an optimal number of foragers to maximize expected payoffs on each turn. In a decentralized collective, individuals share a common belief about the environmental state but make individual choices based on a private threshold. If all individuals choose a threshold that maximizes their own expected return, then all thresholds are identical. Since they share the same belief, everyone makes the same decision: forage when the expected reward is positive, and remain idle when the expected reward is negative. As a result, a collective can quickly detect when an environment switches from a high- to a low-rewarding state. However, when the collective believes the reward state is likely to be low, no individuals forage, and the collective will not detect a switch back to a high-rewarding state. The resulting failure to explore thus leads to starvation even if good times have returned. Conversely, if all individuals adopt low thresholds and forage often, the collective can quickly infer the environmental state but incurs large energetic losses in lean times. Effective collective performance therefore requires a division of labor between a minority that forages even when rewards are unlikely, and a majority that waits for favorable conditions.
With an appropriate distribution of individual decision thresholds, a decentralized collective can achieve the same expected returns as an optimally controlled group. We show this by constructing a threshold distribution that matches the number of foragers deployed under the optimal centralized policy. The optimal strategy of a central planner can be represented by a function, $\pi(g),$ that maps the collective belief about the state of the environment, $g,$ to the number of individuals sent out to forage, $n=\pi(g)$ (Fig. 2A). As the belief that the environment is in a high-reward state increases, the planner commits more foragers [Puterman1994, Topkis1998]. The policy defines a sequence of belief thresholds, $\theta_{n}=\inf\{\theta\in[0,1]:\pi(\theta)\geq n\}$ for $n=1,2,\cdots,N$, where $N$ is the group size, so that multiple thresholds can be the same. Each threshold, $\theta_{n},$ is the smallest belief value at which sending out exactly $n$ foragers is optimal. In particular, the majority of thresholds are the same at the critical belief $\theta_{c}$ (Fig. 3), at which the expected foraging return is zero.
Figure 2: Division of labor enables centralized and decentralized coordination. (A) In decentralized collectives, individuals have different foraging thresholds. Risk-averse “deliberators” share a common threshold, $\theta_{c},$ and forage only when the expected reward is positive. Risk-tolerant “scouts” have lower, heterogeneous thresholds and forage even when expected rewards are negative. (B) Under centralized control, a planner implementing the optimal policy, $\pi(g),$ deploys an increasing number of foragers as the belief increases. An appropriate distribution of individual thresholds allows a decentralized collective to match this allocation (compare panels A and B). (C) The optimal number of risk-tolerant scouts, $N_{s},$ grows sublinearly with population size, NN, so an increasingly small fraction of explorers suffices for efficient foraging. (D) The collective belief, $g^{t},$ tracks environmental fluctuations; scouts continue exploring during lean periods, enabling rapid re-engagement when conditions improve.
The strategy of a decentralized collective is fully determined by a set of private thresholds. Hence, a collective of size $N$ whose individuals have thresholds $\{\theta_{n}\}_{n=1}^{N}$ will send out exactly $n$ foragers when the belief satisfies $\theta_{n}\leq g<\theta_{n+1}$, with the convention $\theta_{N+1}=1$ (Fig. 2B). This choice of thresholds thus results in a number of foragers that equals that dictated by the optimal policy, $\pi(g)$ for any belief $g$. Importantly, this is a consequence of individual decisions based on private thresholds, rather than orders from a central planner. Such a distribution of foraging strategies could arise from evolved or learned differences in risk sensitivity [Seeley1983, Dreller1998, Liang2012], consistent with classic response-threshold models of division of labor [Bonabeau1996, Beshers2001].
Collectives governed by a central decision-maker are inherently vulnerable to the loss or failure of this planner. By contrast, decentralized collectives with heterogeneous individual strategies and are therefore more likely to be evolutionarilare more can be more robust to perturbations, which may promote evolutionary stability [Jandt2013, Couzin2005]. As long as phenotypic differences across the population generate an appropriate distribution of risk tolerances, the collective can achieve optimal foraging performance [Seeley1983, Charbonneau2015]. Structured heterogeneity allows decentralized groups to match the allocations of a central planner, yielding robust coordination and sustained performance in dynamic environments. We next analyze the structure of the optimal population, and ask how ecological pressures shape this distribution of strategies.
3.2 Exploratory scouts and synchronized deliberators
Figure 3: Effects of perturbing the distribution of decision thresholds. (A) Optimal threshold distribution for a collective of $N=100$ individuals. (B-D) Perturbed strategy distributions: (B) desynchronizing the risk-averse majority by spreading their thresholds; (C) synchronizing risk-tolerant individuals at a high threshold; (D) synchronizing scouts at a low threshold. Colors indicate strategy types. (E) A trade-off plot of return and accuracy for the optimal strategy and perturbations shows that returns are maximized at intermediate accuracy (gray dot; panel A), reflecting weakened coordinated exploitation under desynchronization (brown–gray curve; panel B), improved accuracy at increased cost under moderate heterogeneity (pink–gray curve; panel C), and degraded returns under excessive risk tolerance (blue–gray curve; panel D). Returns are normalized by the maximum possible return $\rho_{\max}=(\gamma_{+}-\lambda)/2$, achieved by committing the entire collective only in the high reward state.
We next determine the optimal balance between risk-tolerant individuals who forage under uncertainty and risk-averse individuals who forage only when the expected rewards is positive. Empirical studies show a consistent pattern in many social species: a small fraction of individuals, such as “scouts" in honeybee populations, persistently engages in exploration, while most colony members adopt safer strategies and exploit known resources [Seeley1983, Seeley2011, Beekman2001, Dornhaus2006]. We show that the optimal decentralized strategy yields a similar pattern. To do so we use the critical threshold $\theta_{c}$ and classify individuals with decision thresholds below $\theta_{c}$ are risk-tolerant, and those with thresholds equal to or exceeding $\theta_{c}$ as risk-averse.
In small optimal collectives, all individuals are risk-tolerant, because each foraging outcome is highly informative and has a large impact on the collective belief about the environmental state (Fig. 2C). As group size increases, performance improves, but the optimal division of labor changes. The expected cost of foraging grows linearly with the number of active foragers, whereas additional foraging outcomes provide redundant information and thus become less valuable. A mathematical analysis shows that the optimal number of risk-tolerant individuals grows logarithmically with group size (dashed line in Fig. 2C). This sublinear scaling is robust across a wide range of environmental volatilities, foraging costs, and reward structures (SI Appendix, Fig. S1). Consequently, larger groups comprise an increasing number of risk-averse individuals, and a correspondingly smaller fraction of risk-tolerant explorers that achieve efficient collective performance. This prediction is consistent with empirical and theoretical studies showing that effective animal collectives rely on a small exploratory minority whose size grows more slowly than group size [Couzin2005, Cronin2014]. Details of the asymptotic analysis underlying this result are provided in SI Appendix, section 2.
While the decision thresholds of risk-tolerant individuals are distributed, all risk-averse individuals share a common threshold, $\theta_{c}$. Indeed, the entire collective commits to foraging once the expected reward becomes positive (see SI Appendix and Fig. S2). This synchronization is a consequence of the linear increase of total reward with the number of committed individuals, implying that full commitment is optimal when returns are expected to be favorable. Thus, an optimal collective consists of a heterogeneous minority that maintains environmental awareness though continued sampling, and a synchronized majority that maximizes collective returns through coordinated exploitation once conditions are deemed favorable.
Optimal collective foraging strategies exhibit a characteristic cascade of activity in response to environmental changes (Fig. 2D). When conditions deteriorate, widespread foraging produces few successes, quickly shifting the collective belief toward low expected rewards. Most individuals stop foraging, leaving a small subset of risk-tolerant explorers that continue sampling. When conditions deteriorate severely, even limited exploration ceases to minimize costs. Nevertheless, in the absence of new information, collective belief gradually relaxes toward uncertainty, prompting renewed exploration. If some of these sparse foraging attempts are successful, the belief shifts towards high expected rewards, and more foragers are recruited. Once the collective belief crosses the threshold $\theta_{c}$, coordinated collective foraging is triggered again, and the process repeats. Heterogeneity in decision thresholds allows the population to act as a distributed sensor, enabling accurate tracking of environmental changes while optimizing the costs of exploration.
Risk-tolerant individuals thus act as functional leaders by providing information about the environment, rather than by imposing their choices on others. Their exploratory actions inform the more cautious majority, resulting in a form of passive leadership that emerges from threshold heterogeneity[King2009, Strandburg-Peshkin2018, Jolles2020], rather than social dominance or intentional coordination. This dual-speed system driven by the gradual discoveries of risk-takers coupled with rapid collective withdrawal after mass failures in foraging attempts allows for efficient collective responses without explicit coordination (Fig. 2D and SI Appendix, Fig. S3).
3.3 Uniform strategies degrade collective performance
A division of labor based on diverse risk tolerances is therefore crucial for balancing the costs and benefits of exploration (Fig. 3A). We next compare the performance of such optimal collectives to those with suboptimal strategy compositions. We quantify collective performance using two metrics: long-term average return $\rho$, normalized by the maximum possible return $\rho_{\max}$, and accuracy $\alpha$, defined as the time-averaged proportion of individuals committed to foraging when the environment is in the high-reward state (Fig. 3E).
When all thresholds are high, the population is overly cautious, since even relatively risk-tolerant individuals forage too infrequently to detect environmental improvements. As a result the collective responds slowly when good times return, reducing long-term average reward (Fig. 3C and E). In the extreme case where all individuals require excessive certainty to forage ($\theta_{n}>1/2$ for all $n$), unsuccessful foraging drives the collective belief below $1/2$, resulting in the cessation of foraging. In the absence of new observations, the belief relaxes toward complete uncertainty ($g=1/2$). Since all decision thresholds exceed this value, the collective remains inactive indefinitely.
Conversely, a collective with too many risk-takers fails in the opposite way: eager explorers can maintain high accuracy, but only by sustaining near-continuous sampling, so the marginal information gain of these scouts is outweighted by their high energetic cost, depressing long-term returns (Fig. 3D and E). In contrast, unstructured deviations from a shared risk-averse threshold primarily impair exploitation rather than information acquisition, desynchronizing the cautious majority and preventing coordinated surges of foraging when conditions are favorable (Fig. 3B and E).
Optimal collectives thus exhibit functional specialization, with a synchronized risk-averse majority sharing threshold $\theta_{c}$ responsible for coordinated exploitation, and a heterogeneous risk-tolerant minority sustaining exploration through a broad distribution of thresholds in large groups (Fig. 3 and SI Appendix, Fig. S4). Together, these results reveal a collective exploration–exploitation trade-off: heterogeneity among risk-tolerant individuals sustains exploration under uncertainty, while synchronization of the risk-averse majority maximizes returns once conditions are favorable. We next ask how ecological conditions reshape this optimal division of labor, when heterogeneity is favored, and when it collapses in favor of homogeneous strategies as the environment change from benign to volatile extremes.
3.4 Ecological extremes lead to a collapse in the division of labor
[…]
Group heterogeneity varies non-monotonically with foraging cost, peaking at intermediate values (Fig. 4A and D, SI Appendix, Fig. S5). When costs are low ($\theta_{c}(\lambda)<\epsilon$), resources are easily accessed and all individuals forage, resulting in a homogeneous population of foragers. Conversely, when costs are high ($1-\epsilon<\theta_{c}(\lambda)$), exploration becomes prohibitively expensive, and all individuals refrain from foraging. Between these extremes ($\epsilon\leq\theta_{c}(\lambda)\leq 1-\epsilon$), a maximal variability of risk-preference is attained. In this intermediate regime, moderate increases in foraging cost selectively suppress marginal exploration by cautious individuals, while favoring continued sampling by a smaller risk-tolerant subset, thereby amplifying the diversity of thresholds used (wedge-shaped region of high heterogeneity in Fig. 4D).
[…]
Rapid environmental changes suppress heterogeneity, favoring homogeneous risk aversion over exploratory strategies (Fig. 4C). When environments fluctuate rapidly, maintaining risk-tolerant individuals becomes counterproductive as information gained through exploration quickly becomes obsolete. In this regime, the timescale of environmental changes is shorter than the timescale required for exploration and coordinated response. As a result, fast-changing environments favor conservative strategies that minimize exploration costs, leading to uniform populations of risk-averse individuals. By contrast, when environmental changes are slow and predictable relative to the timescale of exploration and response, information accumulates and remains reliable over long periods, so the group can profit from exploratory foraging in future decisions. Under these conditions, risk-tolerant individuals can repeatedly detect and confirm state changes, stabilizing collective belief and enabling timely collective commitment, favoring heterogeneous strategies.
Discussion
We have shown that leaderless collectives can achieve optimal foraging performance through a division of labor based on heterogeneous risk tolerances. A small fraction of risk-tolerant individuals, whose number increases logarithmically with group size, maintains environmental awareness by exploring even when expected rewards are negative, while a synchronized risk-averse majority maximizes collective gains by committing only when conditions are favorable. This asymmetric structure reflects fundamental principles of collective information gathering: as individuals sample independently, the information value of additional observations diminishes due to redundancy, whereas exploration costs accumulate linearly with the number of scouts. As a result, only a small fraction of explorers is needed for accurate environmental tracking in large groups.
Optimal collective foraging exhibits characteristic decision cascades in response to environmental changes. Minority-driven exploration triggers rapid group-wide recruitment when conditions improve, enabling swift responses to opportunities while limiting costly false alarms. Exploration loses its value under extreme ecological pressures or in rapidly changing environments where information becomes obsolete before it can guide actions. This causes collectives to favor uniform risk aversion over strategic diversity. Hence, heterogeneity in decision thresholds is most pronounced under intermediate foraging costs and environmental contrasts (Fig. 4).
Risk-tolerant individuals function as informative leaders in decentralized groups, guiding collective behavior by maintaining environmental awareness during periods of uncertainty [Strandburg-Peshkin2018, Couzin2005, King2009]. Whether such leadership is persistent or transient depends on its mechanistic origin. Stable phenotypic variation in energetic state or temperament can produce consistent leaders [Biro2008, King2009, Jolles2017], whereas probabilistic decision-making by otherwise homogeneous individuals allows leadership to rotate while preserving near-optimal collective performance [Couzin2011, Conradt2005, Jolles2020] (See SI Appendix, Section 4 for proof).
While comparable divisions of labor can be imposed by centralized or dominance-based control in some systems [Conradt2005, Mech1999], we focused on how informational leadership can emerge in leaderless collectives via heterogeneity in exploration under uncertainty [Couzin2009, Hein2015]. A minority that samples when payoffs are ambiguous can distribute group sensing to regulate majority engagement, consistent with broader views of collective computation and information flow in animal groups [Sumpter2008]. This perspective suggests that if exploratory roles reflect stable phenotypic differences, individuals should show cross-context consistency in risk-taking and influence [Biro2008, Jolles2017, King2018]; if instead exploration arises from stochastic choice among otherwise similar individuals, leadership should rotate without loss of group-level performance [Couzin2011, Mann2018]. Correspondingly, altering exploration costs or environmental persistence is expected to shift the degree and stability of leadership and the strength of division of labor [Lima1990, Dall2005]. These predictions are experimentally accessible in social insects and other tractable systems where ecological timescales and costs can be manipulated and collective responses quantified [Seeley1994, Davidson2016, Dornhaus2006].
A sublinearly increasing number of information leaders appears across diverse collective decision-making systems: informed minorities can effectively guide naive majorities in moving animal groups [Couzin2005]; a small fraction of hasty explorers can optimize group decision efficiency when observed by a large, deliberate majority [Karamched2020]; and a few scouts can reliably inform nest-site choices in social insects [Cronin2014]. Although leadership by risk-tolerant minorities has been studied experimentally in groups of fixed size [Reebs2000, Beekman2006, Dyer2008], how the number of leaders should scale with group size and how this scaling emerges from individual decision rules remains largely unexplored.
The role of heterogeneity in group composition has been extensively studied using threshold-based models of division of labor, which typically assume symmetric, normally distributed response thresholds across individuals [Bonabeau1996, Bonabeau1998, Beshers2001, OSheaWheller2017, Ulrich2021]. However, the structure of threshold distributions itself has been largely overlooked. Skewness alone can generate complex relationships between group composition and collective performance. Division of labor in social insects can emerge from multiple interacting factors, beyond threshold variation [Ulrich2021], yet the contribution of the distribution structure remains unexplored. We hypothesize that natural groups may evolve skewed threshold compositions to optimize collective task performance under uncertainty.
One limitation of our model is the assumption of perfect information sharing: all individuals instantaneously observe collective foraging outcomes. In many natural systems, however, not all potential foragers observe every returning individual, and information is transmitted locally or indirectly [Sumpter2005, Sumpter2008, Karamched2020]. Under such conditions, individuals may rely not only on explicit signals but also on implicit social information inferred from the behavior of others [Mann2018, Mann2020, Suganuma2025]. How partial information sharing and social inference reshape the optimal balance of division of labor remains an open question [Beshers2001]. We hypothesize that the core organizational structure identified here – a minority of explorers along with a synchronized majority – remains robust under imperfect communication, although achieving comparable performance may require a larger exploratory cohort to compensate for information loss.
Our results raise unresolved questions about how finite energetic reserves constrain collective decision-making. While our model implicitly assumes unbounded accumulation of returns, biological systems operate under both lower and upper energetic limits. Near starvation, collectives may be forced into riskier strategies to avoid extinction. How such transitions are distributed across heterogeneous groups remains unclear. If risk heterogeneity reflects hunger-driven phenotypic variation, individuals with depleted reserves may become disproportionately risk-tolerant [Theraulaz1998], amplifying the influence of an informed minority during critical periods [King2009]. Conversely, when storage approaches saturation, continued foraging becomes inefficient, potentially suppressing exploration and diminishing the role of exploratory individuals altogether. Understanding how energetic constraints reshape the emergence, persistence, and influence of informed minorities is therefore essential for linking individual state, division of labor, and collective responsiveness of natural systems.
In human society, the parallel to the risk-tolerant scout is thus not the state-sponsored adventurer like Zheng He or Magellan, but the self-motivated explorer: the garage inventor, the citizen scientist, the open-source builder, the high-risk entrepreneur, and the avant-garde artist. These individuals are driven by curiosity and passion and have lower thresholds for action than others. They often bear substantial personal costs, yet when they succeed, they spark movements that drive collective behavior at scale.
4 Methods
Bellman equation and numerical solution
To determine the optimal collective foraging strategy under a centralized planner, we formulate the problem as a Markov decision process in which the system state is given by the collective belief about the environmental state. The goal is to find the strategy that maximizes the long-term average collective return. Using dynamic programming, we derive the Bellman equation characterizing the optimal policy, $\pi(g),$ that maps belief states $g$ to the number of agents committed to foraging. The derivation of the Bellman equation is provided in SI Appendix, section 1.
We compute the optimal policy numerically by discretizing the belief space on the interval $[0,1]$ using a uniform grid of $200$ points and solving the discretized Bellman equation via fixed-point iteration. Iteration is continued until the value function converges to tolerance $10^{-6}$. The resulting policy serves as a benchmark for collective performance and as a reference for constructing decentralized strategies.
Asymptotic analysis and large deviation theory
To understand how the optimal collective strategy, $\pi(g),$ changes with group size $N$, we perform asymptotic analysis of the Bellman equation in the limit of large $N$. This analysis characterizes the leading-order behavior of the optimal policy and shows that only a sublinear number of agents explore (forage) under unfavorable conditions, i.e. when $g<\theta_{c}$. In our model setup, this scaling is logarithmic in $N$.
Rare events associated with future rewards and belief transitions are characterized using large deviation theory. These estimates allow us to quantify the contribution of the current number of committed agents to long-term returns and to derive bounds on the optimal allocation of exploratory effort for large groups. Technical assumptions and full derivations are presented in SI Appendix, section 2.
Perturbations of the optimal policy
To perturb the optimal policy in Fig. 3, we construct strategies by interpolating between the optimal policy $\pi(g)$ and three extreme reference policies $\pi_{\text{ext}}(g)$ corresponding to uniformly bold, uniformly cautious, and desynchronized cautious strategies. More precisely, the perturbed policy is given by $\pi_{\lambda}(g)=\lfloor(1-\lambda)\pi(g)+\lambda\pi_{\text{ext}}(g)\rfloor$ for $\lambda\in[0,1]$, where $\lfloor z\rfloor$ denotes the largest integer less than or equal to $z$. These interpolations systematically alter the degree of heterogeneity and synchronization in foraging behavior. Performance under perturbed policies is evaluated using the same metrics as for the optimal strategy (return and accuracy).
Model parameters
[…]
Monte Carlo simulations
To compute performance metrics for given model parameters, we employed Monte Carlo simulations of the stochastic foraging process. For each parameter set and (perturbed) policy, we simulate the collective dynamics over long time horizons to estimate the performance metrics. Each data point in Fig. 3 is computed from $10^{3}$ independent realizations, each simulated for $2\times 10^{4}$ time steps.

### v2

@@
-Division of labor enables efficient collective decision-making under uncertainty
+Daring few, patient many: division of labor in decentralized foraging collectives
 Abstract
 How do social animals make effective decisions in the absence of a leader?
-While coordination can improve accuracy, it also delays responses as information propagates through the group.
-In changing environments, these delays can outweigh the benefits of centralized control, making decentralized strategies advantageous in large groups.
-This raises a key question: how can groups implement efficient collective decisions without central coordination?
-We address this question using a model of collective foraging in which individuals choose whether to invest in costly exploration or remain idle, while sharing information and rewards across the group.
-We show that decentralized collectives can match the performance of centrally controlled groups through a division of labor: a small, heterogeneous subset explores even when expected rewards are negative, while a synchronized majority forages only when expected rewards are positive.
-Information redundancy causes the optimal scout number to grow sublinearly with group size, so that larger groups need proportionally fewer explorers.
-The heterogeneity of the group is maximized at intermediate ecological pressures, but optimal groups are homogeneous when costs or environmental contrasts or fluctuations are extreme.
+While coordination can improve accuracy, it also introduces delays as information propagates through the group.
+In changing environments, these delays can outweigh the benefits of globally coordinated decisions, even when local interactions remain tightly organized.
+This raises a key question: how can groups implement efficient collective decision-making without central coordination?
+We address this question using a collective foraging model in which individuals share information and rewards, but each must choose whether to bear the cost of exploring or to remain idle.
+We show that decentralized collectives can match the performance of centrally controlled groups through a division of labor: a small, heterogeneous subset explores even when expected rewards are negative, acquiring information to enable future foraging, while a coordinated majority forages only when expected rewards are positive.
+Information redundancy causes the optimal number of explorers to grow sublinearly with group size, so that larger groups need proportionally fewer explorers.
+The heterogeneity of the group is maximized at intermediate ecological pressures, but optimal groups are homogeneous when costs or fluctuations are extreme.
 Crucially, these group-level policies do not require central coordination, emerging instead from agents following simple threshold-based decision rules.
@@
 This tension raises a fundamental question: How should groups be organized to balance these costs and benefits?
-One solution to this problem is centralized leadership.
-In many wild wolf packs the breeding pair tends to initiate movements and coordinates traveling, resting, and hunting [Mech1999].
+One solution to this problem is centralized leadership, in which a small number of individuals guide collective action.
+In many wild wolf packs, for example, the breeding pair tends to initiate movements and coordinates traveling, resting, and hunting [Mech1999].
 Human history offers parallel examples.
 Admiral Zheng He commanded fleets across the Indian Ocean in the early fifteenth century on imperial orders, spurring Chinese emigration and trade throughout Southeast Asia [Levathes2014].
-A century later, royal sponsorship supported Magellan’s circumnavigation, which opened the first western sea route to the Spice Islands and accelerated global trade and contact between distant regions [Bergreen2003].
-In each case, a designated explorer, acting under a central authority, bore enormous personal risk, while the nations that sent them reaped lasting rewards.
-However, most social animals including honeybees, ants, and fish schools lack fixed leaders and coordinate through local interactions [Pratt2002, Seeley2004, Couzin2005, Sumpter2005, Seeley2011, Hein2015].
-Despite extensive empirical study of decentralized systems [Pratt2002, Sumpter2005, Giardina2008, Gordon2014], it remains unclear how closely decentralized coordination matches centralized control, how performance depends on diversity in individual strategies, and under what conditions heterogeneous strategies offer an advantage in leaderless groups.
-To address these questions we develop and analyze a stochastic foraging model [Puterman1994, Mahadevan1996, Gold2007, Kilpatrick2021, BlumMoyse2025, Dembo2010, Bressloff2021, Barendregt2025] to understand how collectives can achieve efficient exploration.
-We show that well-calibrated collectives exhibit characteristic dynamics: during times of plenty, risk-averse individuals venture out together, amplifying collective gains.
-When resources are scarce, most individuals remain idle.
-Yet, a small subset of risk-tolerant explorers continues foraging even when expected rewards are negative, and can rapidly detect environmental improvements.
-Composition is critical as population dominated by cautious individuals fail to discover new resources, while those dominated by risk-tolerant foragers exhaust shared reserves.
-Indeed, the number of risk-tolerant individuals grows sublinearly with population size, making a balanced division of labor between exploration and exploitation essential for efficient foraging [Seeley2011, Charbonneau2015].
+A century later, royal sponsorship supported Magellan’s circumnavigation, opening the first western sea route to the Spice Islands and accelerated global trade and contact between distant regions [Bergreen2003].
+In each case, a small number of designated explorers, acting under a central authority, bore enormous personal risk, while the societies that sent them reaped lasting rewards.
+However, most social animals, including honeybees, ants, and fish schools, lack fixed leaders and coordinate through local interactions [Pratt2002, Seeley2004, Couzin2005, Sumpter2005, Seeley2011, Hein2015].
+Despite extensive empirical study of decentralized animal groups [Pratt2002, Sumpter2005, Giardina2008, Gordon2014], it remains unclear how closely decentralized coordination can match centralized control, how performance depends on diversity in individual strategies, and under what conditions heterogeneous strategies offer an advantage in leaderless groups.
+To address these questions, we develop and analyze a stochastic foraging model [Puterman1994, Mahadevan1996, Gold2007, Kilpatrick2021, BlumMoyse2025, Dembo2010, Bressloff2021, Barendregt2025] to understand how collectives can achieve efficient exploration.
+We show that well-calibrated collectives exhibit two characteristic behavioral regimes: during times of plenty, risk-averse individuals venture out together, amplifying collective gains.
+When resources are scarce, most individuals refrain from foraging, but a small subset of risk-tolerant explorers continues foraging to rapidly detect environmental improvements.
+Composition is critical as populations dominated by cautious individuals fail to discover new resources, whereas those dominated by risk-tolerant foragers exhaust shared reserves.
+The number of risk-tolerant individuals grows sublinearly with population size, making a balanced division of labor between exploration and exploitation essential for efficient foraging [Seeley2011, Charbonneau2015].
 Remarkably, collectives with a well-calibrated mixture of cautious and bold individuals can perform as well as groups governed by a central decision-maker.
 The most heterogeneous compositions emerge under intermediate ecological pressures, when exploration costs and environmental variability are balanced.
-We thus show how simple individual decision rules lead to optimal collective performance without coordination, and predict when ecological pressures favor heterogeneous versus uniform group composition.
+We thus show how simple individual decision rules lead to optimal collective performance and predict when ecological pressures favor heterogeneous versus uniform group composition.
 Figure 1: Overview of model and setup.
@@
 (D) Under centralized control, a planner decides on the optimal number of foragers to maximize returns.
-(E) Under decentralized control, coordination emerges from individual foraging decisions.
+(E) With decentralized control, coordination emerges from individual foraging decisions.
 2 A Model of Collective Foraging under Uncertainty
-Collective decisions about foraging in large, decentralized collectives face a fundamental challenge: How do groups balance individual risk and collective benefit in dynamically changing environments without central coordination?
-To address this question we consider a minimal model that shows that decentralized decision-making mechanisms can yield optimal collective long-term returns, even when individual expected returns are negative.
+Collective decisions about foraging in large, decentralized collectives face a fundamental challenge: how do groups balance individual risk and collective benefit in changing environments without central coordination?
+To address this question, we consider a minimal model designed to examine whether decentralized decision-making mechanisms can yield optimal collective long-term returns, even when individual expected returns are negative.
 Here we provide an overview of the model; mathematical details can be found in SI Appendix, Section 1.
@@
 Returning foragers are observed by the entire community, so all individuals know how many foraging attempts have been successful (Fig.
-1A).
+1A), an idealization of group-level information flow in social collectives [Sumpter2005, Sumpter2008, Couzin2005].
 Individuals are Bayesian observers [Wald1948, Bogacz2006, VelizCuba2016] and compute the probability that the environmental state is in a high-reward state based on the observed number of successful foragers and knowledge about the environment’s volatility.
@@
 Maximizing returns requires optimizing forager number: too many foragers when conditions are poor wastes effort; too few when conditions are good misses opportunities.
-To establish an upper bound on collective performance, we derive a theoretical benchmark using a centralized planner that optimally determines foraging numbers to maximize long-term collective returns (Fig.
+To establish an upper bound on collective performance, we derive a theoretical benchmark based on a centralized planner that optimally determines foraging effort required to maximize long-term collective returns (Fig.
 1D).
-Under this centralized policy, all individuals follow the directives of a planner.
-This planner infers the environmental state by observing foraging successes to determine the best number of foragers.
+Under this centralized policy, all individuals follow the directives of a planner, which infers the environmental state by observing foraging successes and then prescribes the best number of foragers.
 This optimal policy can be computed using dynamic programming [Bellman1954, Mahadevan1996, Sutton1998] and provides an upper bound on achievable performance.
-We compare this benchmark with outcomes from independent agents who make individual decisions based on shared information but without central coordination.
+We compare this benchmark with outcomes from agents who make individual decisions based on shared information but without central coordination.
 Decisions in the absence of a centralized planner
@@
 3 Results
-We use our model of foraging and belief-updating to ask how a collective translates shared information about the environment into an effective foraging decisions.
-We begin by establishing a normative benchmark defined by the allocation selection of a centralized decision-maker to maximize expected payoff.
-We then show that a leaderless group can achieve optimal performance using heterogeneous individual threshold rules, and analyze how environmental conditions shape structured heterogeneity in an optimal, decentralized collective.
+Using our model of foraging and belief updating, we ask how a collective translates shared environmental information into effective foraging decisions.
+We first establish a normative benchmark based on a centralized decision-maker that maximizes expected payoff.
+We then show that a leaderless group can achieve optimal performance using heterogeneous individual decision thresholds and analyze how environmental conditions shape the optimal threshold distribution in a decentralized collective.
 3.1 Robust collective coordination through structured heterogeneity
@@
 However, when the collective believes the reward state is likely to be low, no individuals forage, and the collective will not detect a switch back to a high-rewarding state.
-The resulting failure to explore thus leads to starvation even if good times have returned.
+The resulting failure to explore thus leads to starvation even when conditions improve.
 Conversely, if all individuals adopt low thresholds and forage often, the collective can quickly infer the environmental state but incurs large energetic losses in lean times.
-Effective collective performance therefore requires a division of labor between a minority that forages even when rewards are unlikely, and a majority that waits for favorable conditions.
+Effective collective performance therefore requires a division of labor between a minority that forages even when rewards are unlikely and a majority that waits for favorable conditions.
 With an appropriate distribution of individual decision thresholds, a decentralized collective can achieve the same expected returns as an optimally controlled group.
-We show this by constructing a threshold distribution that matches the number of foragers deployed under the optimal centralized policy.
+We show this by constructing a threshold distribution that reproduces the number of foragers deployed under the optimal centralized policy.
+We do not suggest that collectives compute or implement a centralized solution; rather, this construction demonstrates that simple local threshold rules can recover centralized allocations, allowing us to analyze how ecological parameters shape the structure and robustness of the resulting strategy.
 The optimal strategy of a central planner can be represented by a function, $\pi(g),$ that maps the collective belief about the state of the environment, $g,$ to the number of individuals sent out to forage, $n=\pi(g)$ (Fig.
@@
 As the belief that the environment is in a high-reward state increases, the planner commits more foragers [Puterman1994, Topkis1998].
-The policy defines a sequence of belief thresholds, $\theta_{n}=\inf\{\theta\in[0,1]:\pi(\theta)\geq n\}$ for $n=1,2,\cdots,N$, where $N$ is the group size, so that multiple thresholds can be the same.
-Each threshold, $\theta_{n},$ is the smallest belief value at which sending out exactly $n$ foragers is optimal.
-In particular, the majority of thresholds are the same at the critical belief $\theta_{c}$ (Fig.
-3), at which the expected foraging return is zero.
+The optimal policy, $\pi(g),$ is non-decreasing in the belief $g$, since higher expected rewards never justify deploying fewer foragers.
+We can therefore define a sequence of belief thresholds, $\theta_{n}=\inf\{\theta\in[0,1]:\pi(\theta)\geq n\}$ for $n=1,2,\cdots,N$, where $N$ is the group size.
+Each threshold, $\theta_{n},$ is the smallest belief value at which sending out at least $n$ foragers is optimal.
+Multiple thresholds may coincide, and in large groups the majority of thresholds equal the critical belief $\theta_{c}$ (Fig.
+3) at which the expected foraging return is zero.
 Figure 2: Division of labor enables centralized and decentralized coordination.
 (A) In decentralized collectives, individuals have different foraging thresholds.
-Risk-averse “deliberators” share a common threshold, $\theta_{c},$ and forage only when the expected reward is positive.
-Risk-tolerant “scouts” have lower, heterogeneous thresholds and forage even when expected rewards are negative.
+Risk-averse individuals share a common threshold, $\theta_{c},$ and forage only when the expected reward is positive.
+Risk-tolerant explorers have lower, heterogeneous thresholds and forage even when the expected reward is negative.
 (B) Under centralized control, a planner implementing the optimal policy, $\pi(g),$ deploys an increasing number of foragers as the belief increases.
 An appropriate distribution of individual thresholds allows a decentralized collective to match this allocation (compare panels A and B).
-(C) The optimal number of risk-tolerant scouts, $N_{s},$ grows sublinearly with population size, NN, so an increasingly small fraction of explorers suffices for efficient foraging.
-(D) The collective belief, $g^{t},$ tracks environmental fluctuations; scouts continue exploring during lean periods, enabling rapid re-engagement when conditions improve.
+(C) The optimal number of risk-tolerant explorers, $N_{s},$ grows sublinearly with population size, NN, so an increasingly small fraction of explorers suffices for efficient foraging.
+(D) The collective belief, $g^{t},$ tracks environmental fluctuations because a subset continues exploring during lean periods, so the collective can rapidly re-engage when conditions improve.
 The strategy of a decentralized collective is fully determined by a set of private thresholds.
@@
 2B).
-This choice of thresholds thus results in a number of foragers that equals that dictated by the optimal policy, $\pi(g)$ for any belief $g$.
-Importantly, this is a consequence of individual decisions based on private thresholds, rather than orders from a central planner.
-Such a distribution of foraging strategies could arise from evolved or learned differences in risk sensitivity [Seeley1983, Dreller1998, Liang2012], consistent with classic response-threshold models of division of labor [Bonabeau1996, Beshers2001].
+This choice of thresholds thus results in a number of foragers equal to that dictated by the optimal policy, $\pi(g)$ for any belief $g$.
+Importantly, individual decisions are based on a shared belief paired with private thresholds, rather than on orders from a central planner.
+The resulting distribution of foraging strategies could arise from evolved or learned differences in risk sensitivity [Seeley1983, Dreller1998, Liang2012], and is consistent with classic response-threshold models of division of labor [Bonabeau1996, Beshers2001].
 Collectives governed by a central decision-maker are inherently vulnerable to the loss or failure of this planner.
-By contrast, decentralized collectives with heterogeneous individual strategies and are therefore more likely to be evolutionarilare more can be more robust to perturbations, which may promote evolutionary stability [Jandt2013, Couzin2005].
+In contrast, decentralized collectives with heterogeneous individual strategies are more robust to decision noise and environmental perturbations, potentially promoting evolutionary stability [Jandt2013, Couzin2005].
 As long as phenotypic differences across the population generate an appropriate distribution of risk tolerances, the collective can achieve optimal foraging performance [Seeley1983, Charbonneau2015].
-Structured heterogeneity allows decentralized groups to match the allocations of a central planner, yielding robust coordination and sustained performance in dynamic environments.
-We next analyze the structure of the optimal population, and ask how ecological pressures shape this distribution of strategies.
-3.2 Exploratory scouts and synchronized deliberators
+Structured heterogeneity thus allows decentralized groups to match the allocations of a central planner while maintaining robust performance in dynamic environments.
+We next analyze the structure of the optimal population and ask how ecological pressures shape this distribution of strategies.
+3.2 Minority exploration and majority synchronization
 Figure 3: Effects of perturbing the distribution of decision thresholds.
 (A) Optimal threshold distribution for a collective of $N=100$ individuals.
-(B-D) Perturbed strategy distributions: (B) desynchronizing the risk-averse majority by spreading their thresholds; (C) synchronizing risk-tolerant individuals at a high threshold; (D) synchronizing scouts at a low threshold.
-Colors indicate strategy types.
-(E) A trade-off plot of return and accuracy for the optimal strategy and perturbations shows that returns are maximized at intermediate accuracy (gray dot; panel A), reflecting weakened coordinated exploitation under desynchronization (brown–gray curve; panel B), improved accuracy at increased cost under moderate heterogeneity (pink–gray curve; panel C), and degraded returns under excessive risk tolerance (blue–gray curve; panel D).
-Returns are normalized by the maximum possible return $\rho_{\max}=(\gamma_{+}-\lambda)/2$, achieved by committing the entire collective only in the high reward state.
-We next determine the optimal balance between risk-tolerant individuals who forage under uncertainty and risk-averse individuals who forage only when the expected rewards is positive.
+(B-D) Perturbed strategy distributions: (B) increasing the spread of thresholds in the risk-averse majority; (C) risk-tolerant individuals sharing a high threshold; (D) risk-tolerant individuals sharing a low threshold.
+Colors indicate the different perturbation types.
+(E) Normalized return vs.
+accuracy for the optimal strategy and its perturbations shows that returns are maximized at intermediate accuracy (gray dot; panel A), reflecting weakened coordinated exploitation under desynchronization (brown–gray curve; panel B), improved accuracy at increased cost under moderate heterogeneity (pink–gray curve; panel C), and degraded returns under excessive risk tolerance (blue–gray curve; panel D).
+Returns are normalized by the maximum possible return $\rho_{\max}$, defined as the return obtained when the entire collective forages only in the high-reward state.
+An optimal collective balances risk-tolerant individuals, who forage under uncertainty, with risk-averse individuals, who forage only when the expected reward is positive.
 Empirical studies show a consistent pattern in many social species: a small fraction of individuals, such as “scouts" in honeybee populations, persistently engages in exploration, while most colony members adopt safer strategies and exploit known resources [Seeley1983, Seeley2011, Beekman2001, Dornhaus2006].
 We show that the optimal decentralized strategy yields a similar pattern.
-To do so we use the critical threshold $\theta_{c}$ and classify individuals with decision thresholds below $\theta_{c}$ are risk-tolerant, and those with thresholds equal to or exceeding $\theta_{c}$ as risk-averse.
+To do so, we use the critical threshold, $\theta_{c},$ and classify individuals with decision thresholds below $\theta_{c}$ as risk-tolerant, and those with thresholds equal to or exceeding $\theta_{c}$ as risk-averse.
 In small optimal collectives, all individuals are risk-tolerant, because each foraging outcome is highly informative and has a large impact on the collective belief about the environmental state (Fig.
@@
 A mathematical analysis shows that the optimal number of risk-tolerant individuals grows logarithmically with group size (dashed line in Fig.
-2C).
+2C), reflecting diminishing informational returns from additional explorers as observations become increasingly redundant.
 This sublinear scaling is robust across a wide range of environmental volatilities, foraging costs, and reward structures (SI Appendix, Fig.
 S1).
-Consequently, larger groups comprise an increasing number of risk-averse individuals, and a correspondingly smaller fraction of risk-tolerant explorers that achieve efficient collective performance.
+Consequently, larger groups comprise a growing majority of risk-averse individuals, while the fraction of risk-tolerant explorers declines.
 This prediction is consistent with empirical and theoretical studies showing that effective animal collectives rely on a small exploratory minority whose size grows more slowly than group size [Couzin2005, Cronin2014].
@@
 S2).
-This synchronization is a consequence of the linear increase of total reward with the number of committed individuals, implying that full commitment is optimal when returns are expected to be favorable.
-Thus, an optimal collective consists of a heterogeneous minority that maintains environmental awareness though continued sampling, and a synchronized majority that maximizes collective returns through coordinated exploitation once conditions are deemed favorable.
+This synchronization of foraging bouts is a consequence of the linear increase of total reward with the number of committed individuals, implying that full commitment is optimal when returns are expected to be favorable.
+Thus, an optimal collective consists of a heterogeneous minority that maintains environmental awareness through continued sampling, and a synchronized majority that maximizes collective returns through coordinated exploitation once conditions are deemed favorable.
 Optimal collective foraging strategies exhibit a characteristic cascade of activity in response to environmental changes (Fig.
@@
 Once the collective belief crosses the threshold $\theta_{c}$, coordinated collective foraging is triggered again, and the process repeats.
-Heterogeneity in decision thresholds allows the population to act as a distributed sensor, enabling accurate tracking of environmental changes while optimizing the costs of exploration.
+Heterogeneity in decision thresholds enables the population to function as a distributed sensor, maintaining sensitivity to environmental changes while balancing the cost of exploration.
 Risk-tolerant individuals thus act as functional leaders by providing information about the environment, rather than by imposing their choices on others.
 Their exploratory actions inform the more cautious majority, resulting in a form of passive leadership that emerges from threshold heterogeneity[King2009, Strandburg-Peshkin2018, Jolles2020], rather than social dominance or intentional coordination.
-This dual-speed system driven by the gradual discoveries of risk-takers coupled with rapid collective withdrawal after mass failures in foraging attempts allows for efficient collective responses without explicit coordination (Fig.
+This dual-speed system, driven by the gradual discoveries of risk-takers coupled with rapid collective withdrawal after mass failures in foraging attempts, allows for efficient collective responses without explicit coordination (Fig.
 2D and SI Appendix, Fig.
@@
 We next compare the performance of such optimal collectives to those with suboptimal strategy compositions.
-We quantify collective performance using two metrics: long-term average return $\rho$, normalized by the maximum possible return $\rho_{\max}$, and accuracy $\alpha$, defined as the time-averaged proportion of individuals committed to foraging when the environment is in the high-reward state (Fig.
+We quantify collective performance using two metrics (Fig.
 3E).
+The first is the long-term average return, $\rho$, normalized by the maximum possible return $\rho_{\max}=(\gamma_{+}-\lambda)/2$ achieved when all individuals forage only when conditions are favorable.
+The second is accuracy, $\alpha$, defined as the time-averaged proportion of individuals committed to foraging when the environment is in the high-reward state.
 When all thresholds are high, the population is overly cautious, since even relatively risk-tolerant individuals forage too infrequently to detect environmental improvements.
-As a result the collective responds slowly when good times return, reducing long-term average reward (Fig.
+As a result, the collective responds slowly when good times return, reducing long-term average reward (Fig.
 3C and E).
-In the extreme case where all individuals require excessive certainty to forage ($\theta_{n}>1/2$ for all $n$), unsuccessful foraging drives the collective belief below $1/2$, resulting in the cessation of foraging.
-In the absence of new observations, the belief relaxes toward complete uncertainty ($g=1/2$).
+In the extreme case where all individuals require excessive certainty to forage ($\theta_{n}>1/2$ for all $n$), unsuccessful foraging drives the collective belief below $1/2$, halting further exploration.
+In the absence of new observations, the belief relaxes back toward uncertainty ($g=1/2$).
 Since all decision thresholds exceed this value, the collective remains inactive indefinitely.
-Conversely, a collective with too many risk-takers fails in the opposite way: eager explorers can maintain high accuracy, but only by sustaining near-continuous sampling, so the marginal information gain of these scouts is outweighted by their high energetic cost, depressing long-term returns (Fig.
+Conversely, a collective with too many risk-takers fails in the opposite way: eager explorers can maintain high accuracy, but only by sustaining near-continuous sampling.
+The marginal information gain of these explorers is outweighed by the energetic cost of foraging, depressing long-term returns (Fig.
 3D and E).
@@
 3B and E).
-Optimal collectives thus exhibit functional specialization, with a synchronized risk-averse majority sharing threshold $\theta_{c}$ responsible for coordinated exploitation, and a heterogeneous risk-tolerant minority sustaining exploration through a broad distribution of thresholds in large groups (Fig.
+Optimal collectives thus exhibit functional specialization, with a synchronized risk-averse majority sharing threshold $\theta_{c}$ responsible for coordinated exploitation, and a heterogeneous risk-tolerant minority that sustains exploration through a broad distribution of thresholds in large groups (Fig.
 3 and SI Appendix, Fig.
@@
 Together, these results reveal a collective exploration–exploitation trade-off: heterogeneity among risk-tolerant individuals sustains exploration under uncertainty, while synchronization of the risk-averse majority maximizes returns once conditions are favorable.
-We next ask how ecological conditions reshape this optimal division of labor, when heterogeneity is favored, and when it collapses in favor of homogeneous strategies as the environment change from benign to volatile extremes.
-3.4 Ecological extremes lead to a collapse in the division of labor
+We next ask how environmental conditions shape the optimal balance between explorers and exploiters, and identify when ecological pressures favor heterogeneous group compositions over uniform ones.
+3.4 Ecological extremes suppress division of labor
 […]
@@
 When environments fluctuate rapidly, maintaining risk-tolerant individuals becomes counterproductive as information gained through exploration quickly becomes obsolete.
-In this regime, the timescale of environmental changes is shorter than the timescale required for exploration and coordinated response.
+In this regime, the timescale of environmental changes is shorter than the time needed to coordinate a response.
 As a result, fast-changing environments favor conservative strategies that minimize exploration costs, leading to uniform populations of risk-averse individuals.
-By contrast, when environmental changes are slow and predictable relative to the timescale of exploration and response, information accumulates and remains reliable over long periods, so the group can profit from exploratory foraging in future decisions.
-Under these conditions, risk-tolerant individuals can repeatedly detect and confirm state changes, stabilizing collective belief and enabling timely collective commitment, favoring heterogeneous strategies.
+In contrast, when environmental changes are slow and predictable relative to the timescale of exploration and response, information accumulates and remains reliable over long periods, so the group can profit from exploratory foraging in future decisions.
+Under these conditions, risk-tolerant individuals can repeatedly detect state changes, stabilizing collective belief and enabling timely collective commitment, favoring heterogeneous strategies.
 Discussion
@@
 A small fraction of risk-tolerant individuals, whose number increases logarithmically with group size, maintains environmental awareness by exploring even when expected rewards are negative, while a synchronized risk-averse majority maximizes collective gains by committing only when conditions are favorable.
-This asymmetric structure reflects fundamental principles of collective information gathering: as individuals sample independently, the information value of additional observations diminishes due to redundancy, whereas exploration costs accumulate linearly with the number of scouts.
-As a result, only a small fraction of explorers is needed for accurate environmental tracking in large groups.
+This asymmetric structure reflects fundamental principles of collective information gathering: as individuals sample independently, the information value of additional observations diminishes due to redundancy, whereas exploration costs accumulate linearly with the number of explorers, so only a small fraction is needed for accurate environmental tracking in large groups.
 Optimal collective foraging exhibits characteristic decision cascades in response to environmental changes.
@@
 Risk-tolerant individuals function as informative leaders in decentralized groups, guiding collective behavior by maintaining environmental awareness during periods of uncertainty [Strandburg-Peshkin2018, Couzin2005, King2009].
-Whether such leadership is persistent or transient depends on its mechanistic origin.
+Whether such leadership is persistent or transient depends on its origin.
 Stable phenotypic variation in energetic state or temperament can produce consistent leaders [Biro2008, King2009, Jolles2017], whereas probabilistic decision-making by otherwise homogeneous individuals allows leadership to rotate while preserving near-optimal collective performance [Couzin2011, Conradt2005, Jolles2020] (See SI Appendix, Section 4 for proof).
-While comparable divisions of labor can be imposed by centralized or dominance-based control in some systems [Conradt2005, Mech1999], we focused on how informational leadership can emerge in leaderless collectives via heterogeneity in exploration under uncertainty [Couzin2009, Hein2015].
+While comparable divisions of labor can be imposed by centralized or dominance-based control in some groups [Conradt2005, Mech1999], we focused on how informational leadership can emerge in leaderless collectives via heterogeneity in exploration preference under uncertainty [Couzin2009, Hein2015].
 A minority that samples when payoffs are ambiguous can distribute group sensing to regulate majority engagement, consistent with broader views of collective computation and information flow in animal groups [Sumpter2008].
-This perspective suggests that if exploratory roles reflect stable phenotypic differences, individuals should show cross-context consistency in risk-taking and influence [Biro2008, Jolles2017, King2018]; if instead exploration arises from stochastic choice among otherwise similar individuals, leadership should rotate without loss of group-level performance [Couzin2011, Mann2018].
+This perspective suggests that if exploratory roles reflect stable phenotypic differences, individuals should show cross-context consistency in risk-taking and influence [Biro2008, Jolles2017, King2018]; if instead exploration is due to stochastic choices among otherwise similar individuals, leadership should rotate without loss of group-level performance [Couzin2011, Mann2018].
 Correspondingly, altering exploration costs or environmental persistence is expected to shift the degree and stability of leadership and the strength of division of labor [Lima1990, Dall2005].
 These predictions are experimentally accessible in social insects and other tractable systems where ecological timescales and costs can be manipulated and collective responses quantified [Seeley1994, Davidson2016, Dornhaus2006].
-A sublinearly increasing number of information leaders appears across diverse collective decision-making systems: informed minorities can effectively guide naive majorities in moving animal groups [Couzin2005]; a small fraction of hasty explorers can optimize group decision efficiency when observed by a large, deliberate majority [Karamched2020]; and a few scouts can reliably inform nest-site choices in social insects [Cronin2014].
+A sublinearly increasing number of information leaders appears in diverse collective decision-making systems: informed minorities can effectively guide naïve majorities in moving animal groups [Couzin2005], a small fraction of hasty explorers can optimize group decision efficiency when observed by a large, deliberate majority [Karamched2020], and a few scouts can reliably inform nest-site choices in social insects [Cronin2014].
 Although leadership by risk-tolerant minorities has been studied experimentally in groups of fixed size [Reebs2000, Beekman2006, Dyer2008], how the number of leaders should scale with group size and how this scaling emerges from individual decision rules remains largely unexplored.
@@
 Division of labor in social insects can emerge from multiple interacting factors, beyond threshold variation [Ulrich2021], yet the contribution of the distribution structure remains unexplored.
-We hypothesize that natural groups may evolve skewed threshold compositions to optimize collective task performance under uncertainty.
-One limitation of our model is the assumption of perfect information sharing: all individuals instantaneously observe collective foraging outcomes.
-In many natural systems, however, not all potential foragers observe every returning individual, and information is transmitted locally or indirectly [Sumpter2005, Sumpter2008, Karamched2020].
+We hypothesize that natural groups may evolve skewed threshold compositions to optimize collective task performance.
+One limitation of our model is the assumption of perfect information sharing as all individuals instantaneously observe collective foraging outcomes.
+In many natural systems not all potential foragers observe every returning individual, and information is transmitted locally or indirectly [Sumpter2005, Sumpter2008, Karamched2020].
+Such local transmission can generate modular organization, in which distinct subpopulations have unequal access to exploratory returns.
+Even when information is available, individuals may differ in perceptual or decision noise (e.g., noisy belief updates or variable thresholds) [Bogacz2006], resulting in heterogeneous beliefs and stochastic actions.
+We expect the qualitative division-of-labor strategy identified here to persist with such noise.
 Under such conditions, individuals may rely not only on explicit signals but also on implicit social information inferred from the behavior of others [Mann2018, Mann2020, Suganuma2025].
-How partial information sharing and social inference reshape the optimal balance of division of labor remains an open question [Beshers2001].
+How partial information sharing and social inference reshape the optimal division of labor remains an open question [Beshers2001].
 We hypothesize that the core organizational structure identified here – a minority of explorers along with a synchronized majority – remains robust under imperfect communication, although achieving comparable performance may require a larger exploratory cohort to compensate for information loss.
-Our results raise unresolved questions about how finite energetic reserves constrain collective decision-making.
+Our results also raise questions about how finite energetic reserves constrain collective decision-making.
 While our model implicitly assumes unbounded accumulation of returns, biological systems operate under both lower and upper energetic limits.
@@
 If risk heterogeneity reflects hunger-driven phenotypic variation, individuals with depleted reserves may become disproportionately risk-tolerant [Theraulaz1998], amplifying the influence of an informed minority during critical periods [King2009].
-Conversely, when storage approaches saturation, continued foraging becomes inefficient, potentially suppressing exploration and diminishing the role of exploratory individuals altogether.
+Conversely, when storage approaches saturation, continued foraging becomes unnecessary, potentially suppressing exploration and diminishing the role of exploratory individuals altogether.
 Understanding how energetic constraints reshape the emergence, persistence, and influence of informed minorities is therefore essential for linking individual state, division of labor, and collective responsiveness of natural systems.
-In human society, the parallel to the risk-tolerant scout is thus not the state-sponsored adventurer like Zheng He or Magellan, but the self-motivated explorer: the garage inventor, the citizen scientist, the open-source builder, the high-risk entrepreneur, and the avant-garde artist.
+In human society, the parallel to the risk-tolerant explorer is thus not the state-sponsored adventurer like Zheng He or Magellan, but the self-motivated innovator: the garage inventor, the citizen scientist, the open-source builder, the high-risk entrepreneur, and the avant-garde artist.
 These individuals are driven by curiosity and passion and have lower thresholds for action than others.
@@
 To understand how the optimal collective strategy, $\pi(g),$ changes with group size $N$, we perform asymptotic analysis of the Bellman equation in the limit of large $N$.
-This analysis characterizes the leading-order behavior of the optimal policy and shows that only a sublinear number of agents explore (forage) under unfavorable conditions, i.e.
-when $g<\theta_{c}$.
+This analysis characterizes the leading-order behavior of the optimal policy and shows that only a sublinear number of agents explore (forage) under unfavorable conditions, i.e., when $g<\theta_{c}$.
 In our model setup, this scaling is logarithmic in $N$.